The temperature of n moles of ideal gas is increased from T to 4T through a process for which pressure P = a T^(-1) ,where 'a' is a constant. The work done by the gas is ?
1
Expert's answer
2020-06-21T20:01:52-0400

Change in temperature is ,"\u0394T=4T\u2212T=3T"

Given, gas obeys the law

"P=aT^{-1}"

Now, from ideal gas law, "PV=nRT"

and solving the above equation we get,

"PV^{\frac{1}{2}}=constant"

Thus, process is poly-tropic.

We, know that molar heat capacity is

"C=C_v+\frac{R}{1-x}"

​Hence,

"C=C_v+2R"

Now, we also know that, heat absorbed is

"ΔQ=nCΔT=ΔU+ΔW=nC_v\Delta T+\Delta W"

Therefore, we get, work done is

"\Delta W=n\Delta T(C-C_v)=6nRT"

Note: "R" is universal gas constant.